OpenStudy (anonymous):

Ok so i am terrible at simplifying. Can someone break this question down and explain it to me? Simplify 5x + 2(x - 3) = -2(x - 1)

OpenStudy (phi):

first, 2 (x-3) means 2 times every thing inside the parens. think of the parens as a "package" and you have two of them. then you have 2 of what ever is inside. that means 2(x-3) means you have 2 x's and 2 -3's or 2 * x + 2*-3

hero (hero):

5x + 2(x - 1 - 2) = -2(x - 1) 5x + 2(x - 1) - 4 = -2(x - 1) 5x + 2(x - 1) + 2(x - 1) - 4 = 0 5x + 4(x - 1) - 4 = 0 5x + 4x - 4 = 4 5x + 4x = 4 + 4 (5 + 4)x = 8 9x = 8 x = 8/9

hero (hero):

You won't be able to understand how to simplify algebraic expressions until you master the distributive rule: a(b + c) = ab + ac